OpenStudy (anonymous):

Write an equation for the line parallel to y = 3 + 2x that passes through the point (3, 2)

OpenStudy (anonymous):

well you know that the slope of your line will be 2 correct from y=mx+b?

OpenStudy (anonymous):

if you are looking for a line parallel to another line, their slopes will be the same

OpenStudy (anonymous):

so you have 2 = your slope

OpenStudy (anonymous):

now use your point slope formula (i think that is what it is) (y-y1)=m(x-x1)

OpenStudy (anonymous):

where y1 = 2 and x1 = 3

OpenStudy (anonymous):

y-2=2(x-3)

OpenStudy (anonymous):

y-2=2x-6 y=2x-4

OpenStudy (anonymous):

so the equation y=2x-4 passes through point (2,3) and is parallel to y= 3+2x

OpenStudy (anonymous):

An equation for the line parallel to "y = 3 + 2x" will have the same slope as "y = 3 + 2x" the slope of "y = 3 + 2x" is 2. so the equation you are looking for will be "y = z + 2x" where z is something you haven't figured out yet. To figure out was z is, plug in the x and y values of the point your equation must pass through (x = 3 and y =2) y = z + 2x plug-in x=3 and y=2 \[2 = z + 2*3\] 2= z + 6 -4 = z so the equation your looking for is "y = -4 + 2x"

OpenStudy (anonymous):

thanks